OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge e^{\theta} (\tan \theta - \theta)\] i have no clue..maybe product rule?

OpenStudy (turingtest):

yeah, product rule

OpenStudy (lgbasallote):

\[\huge e^\theta \tan \theta = e^\theta \theta\] then twice product rule?

OpenStudy (lgbasallote):

that equal is a minus sign of course

OpenStudy (turingtest):

no need to distribute

OpenStudy (lgbasallote):

no?

OpenStudy (lgbasallote):

oh i see

OpenStudy (turingtest):

\[u=e^{\theta}\]\[v=(\tan\theta-\theta)\]

OpenStudy (lgbasallote):

u = \(e^\theta\) v = \((\tan \theta - \theta)\)

OpenStudy (lgbasallote):

yeah what i meant

OpenStudy (turingtest):

ditto!

OpenStudy (lgbasallote):

so it's \[e^\theta (\sec^2 \theta + 1) + e^\theta (\tan \theta -1)?\]

OpenStudy (turingtest):

why the -1 in the second set of parentheses?

OpenStudy (turingtest):

or the +1 in the first...?

OpenStudy (lgbasallote):

oh it's supposed to be theta

OpenStudy (lgbasallote):

and that +1 should be -1

OpenStudy (lgbasallote):

\[e^\theta \tan \theta + e^\theta (\tan \theta - \theta)\]

OpenStudy (lgbasallote):

is that better?

OpenStudy (turingtest):

\[u=e^x\implies u'=e^x\]\[v=\tan x-x\implies v'=\sec^2x-1\]\[(uv)'=u'v+v'u=e^x(\tan x-x)+e^x(\sec^2x-1)\]

OpenStudy (lgbasallote):

so im right right? since sec^2 - 1 = tan^2

OpenStudy (lgbasallote):

oh darn another typo >.<

OpenStudy (lgbasallote):

darn latex

OpenStudy (turingtest):

oh yeah, lol

OpenStudy (turingtest):

now I see what you did

OpenStudy (anonymous):

It's ok I make those too @lgbasallote , Turing is just a living android so he doesn't make syntax errors :-3

OpenStudy (lgbasallote):

so just to be clear \[e^\theta (\tan^2 \theta) + e^\theta (\tan \theta - \theta)?\]

OpenStudy (lgbasallote):

agree @agentx5 darn androids

OpenStudy (turingtest):

@agentx5 yeah right... I have everything copy pasted to erase my errors quickly is all ;)

OpenStudy (turingtest):

yes @lgbasallote

OpenStudy (anonymous):

:-D But as long as it's "first d second + second d first" you're good

OpenStudy (lgbasallote):

thanks! :D

OpenStudy (anonymous):

Oh here's one more for you @lgbasallote : Quotient rule (how I remember it)

OpenStudy (anonymous):

Quotient rule: \[\frac{d}{dt}\frac{high}{low} = \frac{((low)d(high)-(high)d(low)}{(low)^2} \]

OpenStudy (anonymous):

It kind of has a sing-song tone to it, makes it easy to remember

OpenStudy (turingtest):

I like to derive the quotient rule, I almost never even use it use the product rule instead

OpenStudy (anonymous):

Eeek! Yeah I try to avoid making things unnecessarily harder :-D But to each their own technique that works for them, and yours clearly works for you

OpenStudy (turingtest):

let \[\frac uv=(uv^{-1})\]\[(uv^{-1})'=u'v^{-1}-v^{-2}v'u={u'v-v'u\over v^2}\]i.e. product rule+chain rule=quotient rule. but whatever is easiest for you is always best :)